# supplement Analyzing Defect Complex Structures of PL6 in SiC: Insights from First-Principles Screening

### v1

@@ III.1 The Possible PL6 Structure
III.2 Potential Experimental Verification
In the above section, we identified the most promising PL6 candidates, and , through first-principles screening based on ZFS, ZPL, and binding energy. More experimental data are needed to verify the identification. In SiC, Si and C have stable isotopes with non-zero nuclear spins, namely and , with nuclear gyromagnetic ratios of -8.465 MHz/T and 10.705 MHz/T, respectively. The hyperfine interaction between nuclear spin and electron spin can be observed in ESR or ODMR spectra, providing additional signatures for defect structure identification.
The hyperfine tensor consists of an isotropic part (the Fermi contact term) and an anisotropic part , which are expressed as:
A key difference between and is the opposite sign of their gyromagnetic ratios. Consequently, replacing a Si atom site with a C atom leads to a significant change in the sign of the hyperfine tensor at that site. However, the sign of the hyperfine tensor is not directly observable in ESR or ODMR spectra. This stems from the fact that nuclear spin states always encompass both positive and negative components, resulting in redshifted and blueshifted fine-structure splitting pairs relative to the main peak. As the sign is difficult to determine experimentally, we adopt , where is along the c-axis (the principal axis of the tensor), as a characteristic parameter for hyperfine coupling. Furthermore, we analyze the statistical distribution of , which is independent of any specific site within the defect complex. A more detailed discussion can be found in the Supplement Material [23]. The results are presented in Fig. 2.
The statistical distribution of for is almost identical to that of kk (see Fig. LABEL:hyper_stat in Supplement Material [23]) and is omitted here for brevity. The primary difference between and kk is a redshift of approximately 0.5 MHz in the [6, 12] MHz range and a blueshift of approximately 0.7 MHz in the [12, 15] MHz range in the distribution. These shifts are attributed to the hyperfine interaction with nuclei. As outlined in Eqs. (3) and (4), the hyperfine tensor is governed by the spin density distribution, which localizes predominantly around the in the positive -direction. While the defect reverses the sign of the hyperfine interaction at its site (discussed earlier), it also induces a slight perturbation to the spin density at neighboring lattice sites. Since the component is positioned outside the region of significant spin density (as shown in Fig. LABEL:spin-density), the perturbation to the spin density is weak. This results in distribution is nearly identical to that of the isolated kk divacancy. In contrast, the distinct behavior observed for originates primarily from spatial perturbations caused by the proximal defect, rather than the sign reversal of the hyperfine interaction itself.
It is important to note that our analysis focuses on comparing the calculated hyperfine coupling of kk and , which may exhibit systematic deviations from experimental values. However, the relative shift in hyperfine coupling is experimentally observable and provides a critical means to validate our theoretical model.
III.3 Other Potential Spin Defects
| defect | spin | (GHz) | (MHz) | KS gap (eV) | (eV) | identification | |
| hh | 2 | 1.506 | 6.4 | 0.1 | 1.118 | – | PL1 |
| kk | 2 | 1.455 | -4.3 | 0.2 | 1.141 | – | PL2 |
| kh | 2 | 1.395 | -61 | 70.8 | 1.111 | – | PL3 |
| hk | 2 | 1.452 | 29 | 109.9 | 1.137 | – | PL4 |
| 0 | 0.000 | 0.0 | – | 0.555 | -4.771 | – | |
| 0 | 0.000 | 0.0 | – | 0.653 | -4.822 | – | |
| 2 | 1.545 | -26.9 | 0.0 | 1.137 | -0.964 | unknown | |
| 2 | 1.607 | -126.1 | 1.1 | 1.154 | -0.937 | unknown | |
| 2 | 1.411 | -23.3 | 177.5 | 0.951 | -0.810 | red-shifted PL | |
| 2 | 1.470 | -109.9 | 4.3 | 0.971 | -0.754 | red-shifted PL | |
| 2 | 1.287 | 144.7 | 0.8 | 0.899 | -0.387 | red-shifted PL | |
| 2 | 1.338 | 158.9 | 1.6 | 0.880 | -0.363 | red-shifted PL | |
| 2 | 0.981 | -6.9 | 0.2 | 0.603 | -0.339 | red-shifted PL | |
| 2 | 1.306 | 417.5 | 5.9 | 1.021 | -0.338 | undetectable | |
| 2 | 1.049 | 94.5 | 1.6 | 0.798 | -0.328 | red-shifted PL | |
| 2 | 1.702 | -384.4 | 7.6 | 1.038 | -0.327 | undetectable | |
| 2 | 1.347 | 459.9 | 7.0 | 1.001 | -0.321 | undetectable | |
| 2 | 1.425 | -439.8 | 11.2 | 0.664 | -0.315 | undetectable | |
| 2 | 1.769 | -433.2 | 8.5 | 1.034 | -0.309 | undetectable | |
| 2 | 1.487 | 29.7 | 0.8 | 1.040 | -0.253 | unknown | |
| 2 | 1.403 | 95.6 | 1.7 | 0.900 | -0.241 | red-shifted PL | |
| 2 | 1.544 | -27.4 | 1.0 | 1.041 | -0.238 | unknown | |
| 2 | 1.478 | -4.3 | 0.3 | 1.176 | -0.204 | PL6 | |
| 2 | 1.523 | -36.7 | 0.6 | 1.138 | -0.183 | unknown | |
| 2 | 1.460 | -43.8 | 0.2 | 1.153 | -0.159 | hk-like | |
| 2 | 1.527 | -14.8 | 179.5 | 1.111 | -0.098 | hh mixed | |
| 2 | 1.319 | -39.1 | 0.6 | 1.031 | -0.092 | unknown | |
| 2 | 1.554 | 0.3 | 0.3 | 1.207 | -0.083 | PL6 | |
| 2 | 1.387 | -32.1 | 0.7 | 1.038 | -0.082 | unknown | |
| 2 | 1.258 | 60.3 | 1.4 | 0.839 | -0.066 | red-shifted PL | |
| 2 | 1.491 | -35.4 | 1.5 | 1.098 | -0.058 | unknown | |
| 2 | 1.486 | -44.5 | 0.6 | 1.001 | -0.047 | red-shifted PL | |
| 2 | 1.484 | -23.3 | 0.2 | 0.903 | -0.041 | red-shifted PL | |
| 2 | 1.453 | 46.1 | 1.3 | 1.124 | -0.037 | hk-like | |
| 2 | 1.477 | 18.7 | 0.6 | 1.151 | -0.035 | hk-like | |
| 2 | 1.458 | -36.8 | 0.1 | 0.900 | -0.029 | red-shifted PL | |
| 2 | 1.464 | 1.0 | 179.8 | 1.149 | -0.024 | kk mixed | |
| 2 | 1.466 | -21.7 | 0.4 | 0.885 | -0.024 | red-shifted PL | |
| 2 | 1.469 | -16.3 | 0.3 | 1.158 | -0.020 | kk mixed | |
| 2 | 1.473 | 9.7 | 0.1 | 1.155 | -0.014 | kk mixed | |
| 2 | 1.518 | -10.0 | 0.1 | 1.127 | -0.014 | hh mixed | |
| 2 | 1.506 | -9.6 | 0.6 | 1.072 | -0.013 | hh mixed | |
| 2 | 1.559 | -13.2 | 0.6 | 1.109 | -0.010 | hh mixed |
In previous sections, we used experimental data to screen the possible PL6 candidates and obtain the most likely candidates. However, there are also many other defects considered in our calculation that serve as good spin defects but are not PL6 candidates. Table 4 lists all spin defects with binding energies at or below -0.01 eV, which are potentially stable, though formation dynamics were not considered. The properties of divacancies primarily originate from the C dangling bonds formed by the within the divacancy. Introducing a nearest-neighbor antisite disrupts these C dangling bonds, resulting in an ground state, as observed in the and configurations. Electron pairing in these defects enhances their stability relative to others. Besides, most -introduced defect complexes exhibit a significant reduction in crystal field splitting, leading to a lower-energy PL (less than 1.0 eV) or complete PL quenching, rendering them absent in PL spectra. Certain defects also display large (300-400 MHz), whose resonances are too widely separated in ESR or ODMR spectra for unambiguous identification. Moreover, for distant antisites, weak interactions with the central divacancy result in small perturbation, making them hard to distinguish from their parent divacancies. Notably, many defects labeled as “kh-like” exhibit spin properties similar to the hk divacancy while retaining ZPL energies near their original values. However, the principal axis of the remains approximately aligned with the c-axis, a feature resolvable through angle-resolved EPR experiments.
Among numerous stable defect complexes identified, only a limited subset demonstrates characteristics worthy of investigation as potential spin centers with distinct properties. Apart from the plausible PL6 candidates, and , eight additional stable defect complexes, highlighted in bold within Table 4, emerge as effective spin centers. Notably, the and exhibit similar structural compositions and comparable properties. Although their KS level gaps are narrower than those of PL1-4, resulting in lower ZPL energies, their and parameters remain comparable to those of PL1-4. A distinct pair, kk+ and hh+, demonstrates reduced ZPL energies alongside significantly smaller values. The most stable spin defect complexes, kk+ and hh+, exhibit ZPL energies within the PL1-4 spectral range. While kk+ shares similarities with the hk defect, it features an enhanced parameter. The hh+ complex displays an exceptionally large value and a parameter exceeding those of PL1-4 while simultaneously demonstrating superior thermodynamic stability based on binding energy analysis. Additional notable configurations include hh+(with elevated ZPL energy, increased magnitude, and higher value relative to the hh defect), and hh+(reduced ZPL energy while maintaining hk-like defect characteristics).
We have not considered exploring the transition profiles [31, 32, 33] of the PL6 candidates and other spin defects, which could serve as additional criteria. The dynamics of the formation of the defect complex could also be explored to further judge the stability of the defect complex and guide the synthesis of the defect complex in experiments.
@@ IV Conclusion
In conclusion, we have identified the most possible PL6 candidates, and , by the first-principles screening on ZFS, ZPL, and binding energy. The antisite defect is the most probable antisite defect in the SiC crystal, and the symmetry of the defect complex is consistent with the experimental hypothesis. The experimental verification of the candidates could be conducted by the ESR or ODMR spectra, focusing on the hyperfine interaction between the nuclear spin and the electron spin. We also found eight other stable defect complexes that could serve as effective spin centers. Future work could include exploring the dynamics of the formation of the defect complex and further experimental verification of the candidates.
@@ Acknowledgements.
The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of University of Science and Technology of China. The vaspkit code [34] was used to generate calculation files and analyze the results. Vesta [35] was used to visualize the structure.References
@@
- Falk et al. [2013] A. L. Falk, B. B. Buckley, G. Calusine, W. F. Koehl, V. V. Dobrovitski, A. Politi, C. A. Zorman, P. X. L. Feng, and D. D. Awschalom, Nat. Commun. 4, 1819 (2013).
@@
- Luo et al. [2023] Q.-Y. Luo, Q. Li, J.-F. Wang, P.-J. Guo, W.-X. Lin, S. Zhao, Q.-C. Hu, Z.-Q. Zhu, J.-S. Xu, C.-F. Li, and G.-C. Guo, Front. Phys. 11, 10.3389/fphy.2023.1270602 (2023).
- Chen et al. [2025] Y. Chen, Q. Zhang, M. Liu, J. Liu, J. Zhou, P. Yu, S. Lin, Y. Teng, W. Yu, Y. Wang, C. Duan, F. Shi, and J. Du, (2025), under review.
@@
- Jin et al. [2021] Y. Jin, M. Govoni, G. Wolfowicz, S. E. Sullivan, F. J. Heremans, D. D. Awschalom, and G. Galli, Phys. Rev. Mater. 5, 084603 (2021).
- Razinkovas et al. [2021] L. Razinkovas, M. Maciaszek, F. Reinhard, M. W. Doherty, and A. Alkauskas, Phys. Rev. B 104, 235301 (2021).
- Quan and Duan [2025] Q. Quan and C.-K. Duan, Phys. Rev. B 111, 075156 (2025).

### v2

@@ III.1 Screening Potential PL6 Structures in DV-antisite Configurations
III.2 Further Validation Using the hyperfine parameter
The previous screening identified and as key candidates for PL6 within the DV-antisite framework. Concurrently, the c-axis-oriented OV(hh) and OV(kk) structures, previously proposed as potential divacancy-like centers [18], also satisfy our initial screening criteria (as shown in Table 2) and thus warrant rigorous comparison. To further refine the identification of PL6 and critically evaluate all these leading candidates, we now analyze their predicted hyperfine interaction signatures. As established in our methodology (Sec. II.4), our analysis focuses on the principal component, , to allow for the most direct and rigorous comparison with available ODMR experimental data.
To initially visualize and compare the overall hyperfine landscape presented by different candidate structures, we examine the statistical distribution of these calculated values for all relevant and nuclei, as depicted in Fig. 3. This global representation offers an intuitive, qualitative assessment of similarities and differences in the density of hyperfine couplings across various frequency ranges. Fig. 1(a) and Fig. 1(b), which specifically depicts a kk divacancy, illustrates the typical atomic environment and key sites generally relevant for hyperfine interactions in various c-axis-oriented divacancies and related defects in 4H-SiC. The calculated distributions in Fig. 3 generally exhibit characteristic regions: weakly coupled distant nuclei ([1, 4] MHz), at sites ([5, 8] MHz), and more strongly coupled at ([9, 12] MHz) and ([12, 15] MHz) sites [36], with a very strong signal from at the site around 80 MHz (omitted for clarity). While these overall distributions provide a valuable qualitative first look, a more quantitative and decisive comparison with available experimental data necessitates focusing on specific, well-characterized hyperfine interactions. Therefore, our subsequent detailed analysis and comparison with experimental findings from Ref. [35] will concentrate on the relative shifts of the well-characterized and sites.
Before proceeding to a detailed comparison of specific defect models, it is crucial to establish the basis of our validation against experimental hyperfine data. DFT calculations of hyperfine parameters can exhibit systematic deviations from experimental absolute values due to approximations in the exchange-correlation functional, pseudopotentials, and other computational factors [37]. However, relative shifts in hyperfine coupling strengths—such as the change observed when a nearby defect perturbs a divacancy, or differences between distinct defect configurations—are generally more robust and less susceptible to these systematic errors. The proven robustness of this comparative method makes it a reliable tool for validating the systems under investigation. Therefore, our primary validation strategy will focus on comparing these calculated relative shifts with experimentally observed trends, rather than on achieving precise agreement in absolute coupling constants. This approach allows for a more reliable assessment of which candidate structure best reproduces the experimentally characterized hyperfine signatures of PL6.
With this validation approach in mind, initial examination reveals that the distribution for the complex is nearly indistinguishable from that of the unperturbed kk divacancy (see Fig. LABEL:hyper_stat in the Supplemental Material [26](a)), making it an improbable source for the distinct hyperfine signature of PL6. More significantly, the candidate, while initially promising based on other metrics, shows a notable discrepancy when its hyperfine predictions are compared to experiment. Specifically, our calculations for this complex predict a redshift of the peak (0.5 MHz) but a blueshift of the peak (0.7 MHz) relative to the kk divacancy. This pattern is inconsistent with the experimental findings for PL6, which report redshifts for both the (0.4 MHz) and (0.7 MHz) sites when compared to the kk reference [35]. Turning to the OV candidates, the OV(kk) complex can also be dismissed. Our calculations predict it would induce blueshifts at both sites(0.5 MHz for and 0.4 MHz for ) relative to the kk divacancy, again conflicting with the experimental shifts of PL6. In striking contrast, the OV(hh) model exhibits excellent agreement with experimental trends. Our calculations for OV(hh) predict redshifts of 0.4 MHz for and 0.6 MHz for sites relative to the kk divacancy. Furthermore, when comparing to the hh divacancy as another reference, PL6 experimentally exhibits blueshifted and hyperfine couplings; the OV(hh) model successfully reproduces this, yielding calculated blueshifts of 0.5 MHz() and 0.2 MHz (). This consistent agreement in relative hyperfine shifts across multiple reference points and specific sites is compelling. Thus, the hyperfine analysis, grounded in the comparison of these robust relative changes, provides strong evidence identifying OV(hh) as the PL6 center.
III.3 Discussion on the PL5 center
Our preceding analysis, particularly the hyperfine validation comparing potential DV-antisite structures against OV models, strongly indicates that OV(hh) is the PL6 center. An OV(hk) origin for the PL5 center was proposed in Ref. [18], based on its calculated ZFS value, ZPL, and formation energy. However, for a basal defect like hk- (and by extension, OV(hk)), the nonzero ZFS value is also a critical parameter for identification, which has not been explicitly discussed for OV(hk) previously. We therefore calculated the ZFS value for OV(hk) and found its magnitude to be 96.9 MHz. Similarly, an value of 150 MHz has been reported by another independent calculation [38]. Compared to the experimental value for PL5, which is approximately only 16 MHz [6, 39], the significantly larger calculated values (96.9 MHz and 150 MHz) represent a pronounced discrepancy that is unlikely to stem from computational inaccuracies alone. Therefore, our findings do not support the OV(hk) configuration as the origin of the PL5 center.
These findings have notable implications, as PL5 and PL6 are often assumed to share a common origin due to their closely matched ZPL energies [6, 39, 8], stability in charge states [9], and qualitatively similar relative concentrations [17] observed experimentally. However, our results indicate a divergence: while OV(hh) emerges as the most probable candidate for PL6, the PL5 center cannot be straightforwardly attributed to a simple OV(hk) structure. However, it remains possible that PL5 originates from a more intricate configuration beyond the model proposed in Ref. [17].
The potential distinction in origin between PL5 and PL6 can be considered in at least two ways:
Both PL5 and PL6 are fundamentally related to oxygen-vacancy, but PL5 involves a more complex structural variation than the simple OV(hk), thereby maintaining a ‘common family’ association.
PL5 and PL6 arise from fundamentally different sources: PL6 is OV(hh)-based, whereas PL5 may originate from an entirely different defect type, such as a divacancy-antisite complex ().
Distinguishing between these two scenarios is essential. Future experimental investigations, including controlled oxygen doping and subsequent analysis of PL5 and PL6 concentrations and properties, could yield valuable insights. Additionally, theoretical studies should extend the search to encompass more intricate oxygen-related defect configurations beyond simple structures for PL5, complemented by experiment explorations.
@@ IV Conclusion
In conclusion, we have performed a systematic first-principles investigation towards elucidating the microscopic origin of the PL6 center in 4H-SiC. Our study initially examined the DV-antisite hypothesis, screening divacancy-antisite () configurations based on zero-field splitting (ZFS), Kohn-Sham (KS) gap, and binding energy. This analysis identified two promising structures, specifically and , as leading candidates within this framework, both exhibiting the requisite symmetry consistent with experimental observations.
To further validate these candidates, we performed a critical comparison of their predicted hyperfine interaction signatures of the most promising DV-antisite candidates with experimental data for PL6 and with those calculated for the alternative OV models. This comparative analysis demonstrated that the OV(hh) structure accurately reproduces the experimental hyperfine signatures, identifying it as the origin of PL6. Additionally, our re-evaluation of the proposed OV(hk) structure for PL5 identified significant discrepancies between its calculated ZFS parameter and experimental observations, raising doubt about its viability as the PL5 center. This indicates that PL5 and PL6 have distinct microscopic origins.
These findings underscore the necessity of integrating multiple theoretical metrics with diverse experimental signatures to achieve reliable defect identification. Future targeted ESR/ODMR experiments, specifically designed to probe the distinct hyperfine signatures of OV(hh) versus complexes, will be crucial for definitively confirming the PL6 structure. Furthermore, our broader screening has identified additional divacancy-antisite complexes that, while not corresponding to PL6, may represent other stable spin defects in SiC worthy of future investigation.
@@ Acknowledgements.
This work was supported by Innovation Program for Quantum Science and Technology (Grant Nos. 2021ZD0302200 and 2021ZD0303204), the National Natural Science Foundation of China (Grant Nos. 12474242 and 12274396), the Anhui Initiative in Quantum Information Technologies (Grant No. AHY050000), the CAS Project for Young Scientists in Basic Research, the Fundamental Research Funds for the Central Universities, and the China Postdoctoral Science Foundation (Grant Nos. 2021M703110 and 2022T150631). The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of University of Science and Technology of China. The vaspkit code [40] was used to generate calculation files and analyze the results. Vesta [41] was used to visualize the structure.References
@@
- Falk et al. [2013] A. L. Falk, B. B. Buckley, G. Calusine, W. F. Koehl, V. V. Dobrovitski, A. Politi, C. A. Zorman, P. X.-L. Feng, and D. D. Awschalom, Nat. Commun. 4, 1819 (2013).
@@
- Shafizadeh et al. [2025] D. Shafizadeh, N. T. Son, I. A. Abrikosov, and I. G. Ivanov, Phys. Rev. B 111, 165201 (2025).
- Li et al. [2025] J. Li, H. Qi, F. Zhou, Y. Song, N. Xu, B. Hong, H. Chen, Y. Dong, and X. Wang, arXiv:2503.22270 (2025).
- Luo et al. [2023] Q.-Y. Luo, Q. Li, J.-F. Wang, P.-J. Guo, W.-X. Lin, S. Zhao, Q.-C. Hu, Z.-Q. Zhu, J.-S. Xu, C.-F. Li, and G.-C. Guo, Front. Phys. 11, 1270602 (2023).
- Chen et al. [2025] Y. Chen, Q. Zhang, M. Liu, J. Liu, J. Zhou, P. Yu, S. Lin, Y. Teng, W. Yu, Y. Wang, et al., arXiv:2504.07558 (2025).
- Bai et al. [2025] R. Bai, D. Liu, S. Chen, and Y.-N. Wu, Phys. Rev. B 111, 014106 (2025).
@@
- Murzakhanov et al. [2021] F. F. Murzakhanov, B. V. Yavkin, G. V. Mamin, S. B. Orlinskii, H. J. von Bardeleben, T. Biktagirov, U. Gerstmann, and V. A. Soltamov, Phys. Rev. B 103, 245203 (2021).
- Falk et al. [2015] A. L. Falk, P. V. Klimov, V. Ivády, K. Szász, D. J. Christle, W. F. Koehl, A. Gali, and D. D. Awschalom, Phys. Rev. Lett. 114, 247603 (2015).
- Son et al. [2006] N. T. Son, P. Carlsson, J. ul Hassan, E. Janzén, T. Umeda, J. Isoya, A. Gali, M. Bockstedte, N. Morishita, T. Ohshima, and H. Itoh, Phys. Rev. Lett. 96, 055501 (2006).
- Ivády et al. [2018] V. Ivády, I. A. Abrikosov, and A. Gali, npj Comput. Mater. 4, 76 (2018).
- Kobayashi et al. [2023] T. Kobayashi, T. Shimura, and H. Watanabe, J. Appl. Phys. 134, 145701 (2023).
- Li et al. [2021] Q. Li, J.-F. Wang, F.-F. Yan, J.-Y. Zhou, H.-F. Wang, H. Liu, L.-P. Guo, X. Zhou, A. Gali, Z.-H. Liu, et al., Natl. Sci. Rev. 9, nwab122 (2021).